How do you solve 3(c-2)/2 = 2(2c+3)/5 ?
Thank you

Answer by unlockmath(1688)

Hello,
First let's get rid of the fractions by multiplying both sides by 10 to give us:
15(c-2) = 4(2c+3) This can be rewritten as:
15c-30=8c+12 Add 30 to both sides to be:
15c=8c+42 Now subtract 8c from both sides to be:
7c=42 Divide both sides by 7 will result in:
c=6
Make sense?
RJ
Check out a new book at:
www.math-unlock.com